OpenStudy (anonymous):

simplify - cos(π−θ)+sin(π/2 +θ)

OpenStudy (mertsj):

cos(theta)+cos(theta)=2cos(theta)

OpenStudy (mertsj):

\[-\cos(\pi-\theta)=\cos \theta\]

OpenStudy (mertsj):

Also \[\sin(\frac{\pi}{2} +\theta)=\cos \theta\]

OpenStudy (anonymous):

so the answer would be zero?

OpenStudy (mertsj):

Depends on what theta is

OpenStudy (anonymous):

no number to it

OpenStudy (mertsj):

The answer is 2cos(theta) Did you substitute the values in?

OpenStudy (dumbcow):

is there supposed to be a neg sign in front of the cos? "simplify -cos(pi-A)"

OpenStudy (anonymous):

i got it thank you

OpenStudy (anonymous):

can you help me with this one

OpenStudy (anonymous):

4sin6α/cos^2 3α−sin^2 3α